# On the localized and delocalized modes in kink-antikink interactions: a toy model

### v1

@@ III Stability equation
for even and odd delocalized modes, respectively. Once the frequency $\omega$ is computed, the analytical expression of the corresponding bound mode can be obtained as described in Appendix B. In Fig. 2(b), the delocalized modes of the presented potential for the fixed values $A^{2}=3$, $C^{2}=6$ are also shown. Moreover, the values of $\omega^{2}$ as a function of $D/2$ for several values of $A^{2}$ and $C^{2}$ are shown in Figs. 7, 9 and 12.
@@ IV.1 $(A^{2},C^{2})=(8/3,8/3)$
IV.2 $(A^{2},C^{2})=(1.5,7)$
Now, we can consider a case without a localized mode but containing delocalized modes, which can be achieved by fixing $A^{2}=1.5$ and $C^{2}=7$. Resonant behavior via delocalized modes was first described in the $\phi^{6}$ model Dorey2011Kink . Here, we construct a similar scenario where resonance windows can occur via the excitation of delocalized modes. In Fig. 6, we show the field at the collision center as a function of time and initial velocity. In both cases, resonance windows are observed, although they are narrower in the kink-antikink scenario as before. Our toy model not only replicates the resonant structure observed in antikink-kink collisions, similar to the $\phi^{6}$ model, but also reveals a resonant structure in kink-antikink collisions, provided that the meson masses of the two vacua are swapped.
@@
The resonant frequency obtained from the numerical simulations are $\omega_{K\bar{K}}=1.371$ and $\omega_{\bar{K}K}=1.381$. To evaluate whether such values agree with the resonant energy exchange mechanism via the excitation of delocalized modes, we present the analytical delocalized modes frequencies as a function of the separation between the kinks in Fig. 7. As vertical lines, we plot the maximum distance between the kinks in the numerical simulations of the first four bounce windows. The horizontal lines mark the measured frequency. The lowest is the only delocalized mode compatible with the measured frequency in the range delimited by the vertical lines. Therefore, our results agree with the current understanding of the phenomenon.

### v2

@@ III Stability equation
for even and odd delocalized modes, respectively. Once the frequency $\omega$ is computed, the analytical expression of the corresponding bound mode can be obtained as described in Appendix B. In Fig. 2(b), the delocalized modes of the presented potential for the fixed values $A^{2}=3$, $C^{2}=6$ are also shown. Moreover, the values of $\omega^{2}$ as a function of $D/2$ for several values of $A^{2}$ and $C^{2}$ are shown in Figs. 7, 10 and 13.
@@ IV.1 $(A^{2},C^{2})=(8/3,8/3)$
IV.2 $(A^{2},C^{2})=(1.5,7)$ and $(A^{2},C^{2})=(1.1,5.5)$
Now, we can consider two cases, both without a localized mode but containing delocalized modes. They can be achieved by fixing $A^{2}=1.5$ and $C^{2}=7$ for the first scenario and $A^{2}=1.1$ and $C^{2}=5.5$ for the second one. Here, we construct resonance windows via the excitation of delocalized modes, first described in Dorey2011Kink for the $\phi^{6}$ model.
In Fig. 6, we show the field at the collision center as a function of time and initial velocity for $A^{2}=1.5$ and $C^{2}=7$. In both cases, resonance windows are observed, although narrower in the kink-antikink scenario as before. Our toy model not only exhibits the resonant structure expected for antikink-kink collisions but also reveals a resonant structure in kink-antikink collisions, provided that the meson masses of the two vacua are swapped. The resonant frequency obtained from the numerical simulations are $\omega_{K\bar{K}}=1.371$ and $\omega_{\bar{K}K}=1.381$. To evaluate whether such values agree with the resonant energy exchange mechanism via the excitation of delocalized modes, we present the analytical delocalized modes frequencies as a function of the separation between the kinks in Fig. 7. As vertical lines, we plot the maximum distance between the kinks in the numerical simulations of the first four bounce windows. The horizontal lines mark the measured frequency. The lowest is the only delocalized mode compatible with the measured frequency in the range delimited by the vertical lines. Therefore, our results agree with the current understanding of the phenomenon.
A notable feature of the $\phi^{6}$ model is the appearance of not only resonance windows, but also false ones. By adjusting the parameters $A$ and $C$, we can investigate if our toy model can reveal the same feature. In Fig. 8, considering $A^{2}=1.1$ and $C^{2}=5.5$, a mixed structure of false and true windows appears. In Dorey2011Kink , the authors argued that false resonance windows are the result of competition between the lowest delocalized mode and higher frequency ones. Here, we can also associate the appearance of false windows with the change in the spectrum from the first scenario, $A^{2}=1.5$ and $C^{2}=7$.
@@
Figure 8: Field at the collision center as a function of time and $v_{0}$. Antikink-kink collisions are depicted in (a) and the zoom in (b). Parameters are $A^{2}=1.1$ and $C^{2}=5.5$.